OpenStudy (anonymous):

What are the coordinates of the midpoint of the line segment whose endpoints are (2a,3b) and (6a,b)? A) (4a,2b) B) (3a,2b) C) (2a,b) D) (2b,4a)

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

I got 4a,1.5b

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

4a is correct (3b+b)/2 = 4b/2 = 2b

OpenStudy (anonymous):

So, A is right?

hartnn (hartnn):

yes! :)

OpenStudy (anonymous):

Yay!